# SPARTUN3D: Situated Spatial Understanding of 3D World in Large Language Models

### v1

@@ 5.2 Experimental Results
Spartun3D Benchmark
We evaluate the performance of both the LEO model and Spartun3D-LLM after fine-tuning them on our proposed Spartun3D dataset. The fine-tuned LEO model is referred to as LEO+Spartun3D. Table 2 and Table 4 show the experimental results on Situated Captioning and Situated QA tasks, respectively. We experiment with two different LLM backbones: Opt1.3B and Vicuna7B. Our experiments show that Spartun3D-LLM consistently outperforms LEO+Spartun3D across all question types (around $2\%-3\%$ across all metrics), regardless of the LLM backbone used, indicating the effectiveness of our explicit alignment module. We observe that initializing our model with LEO pre-trained weights improves performance. Notably, without fine-tuning, LEO performs reasonably well on attribute and relation questions in a zero-shot setting but struggles with other situated tasks.
SQA3D Performance.
We further evaluate our method on the SQA3D dataset, whose scenes are derived from ScanNet (Dai et al., 2017). Their scenes differ from those in Spartun3D, which are sourced from 3RScan. We experiment with two settings: zero-shot and fine-tuning. In the zero-shot setting, we re-trained LEO exclusively on a dataset constructed from 3RScan to ensure a fair comparison with our method. As shown in Table 3, LEO performs poorly on SQA3D in the zero-shot setting, suggesting its limitations in learning situated understanding from its original dataset. In contrast, LEO trained on Spartun3D shows significant improvement, demonstrating the effectiveness of our dataset and the generalization ability of our model. Furthermore, in the fine-tuning setting, Spartun3D-LLM continues to outperform LEO across all metrics. Table 9 in the Appendix provides a breakdown of the fine-tuned performance across various question types, where Spartun3D-LLM shows consistent improvement across nearly all question types.
@@
Navigation Performance.
To demonstrate the effectiveness of our approach on downstream embodied tasks, we evaluate it on the object navigation tasks. Specifically, we randomly select $5$ scenes that contain around $1000$ examples from the MP3D ObjNav dataset. In this task, we additionally input 2D ego-centric images to both LEO and Spartun3D-LLM for comparison. There are four types of navigation actions: turn left, turn right, move forward, and stop. We evaluate whether the model generates correct action at each step. We conduct the experiment in a zero-shot setting, and Table 5 shows the accuracy of the model’s performance. The baseline model, LEO, struggles to generate the required action-related text to guide navigation steps without fine-tuning specifically for navigation tasks. In contrast, our model demonstrates strong transferability to generate correct actions. Fig 7 (e) showcases a qualitative example, illustrating how our model effectively generates accurate navigation actions without task-specific fine-tuning.
@@
Figure 7: Qualitative Examples. (a), (c), (c) situated qa examples of object attribute and relation, object affordance, and situated planning, respectively. (d) situated captioning example. (e) navigation example in a zero-shot setting. Our model generates both actions and descriptions of surrounding objects while navigating.
@@ 5.3 Ablation Study and Extra Analysis
Explicit Alignment Enhances General Spatial Understanding.
We evaluate the effectiveness of our proposed situated spatial alignment module on general scene understanding tasks, such as
@@
Scan2Cap (Chen et al., 2021) and ScanQA (Azuma et al., 2022). In line with our approach for situated tasks, we construct textual descriptions for each object based on its attributes and spatial relations to others from a top-view perspective. As shown in Tab. 6, by incorporating the explicit spatial alignment module, our model shows better results, indicating that our proposed alignment module not only improves situated understanding but also enhances general 3D scene understanding.
Improved Situated Understanding.
To analyze the model’s situated understanding ability further, we visualize the distribution of model responses generated for questions requiring strong spatial understanding from SQA3D. Specifically, we extract questions starting with “which direction”. Fig. 9 illustrates the distribution of generated “directions”, including “left”, “right”, “forward” and “backward”. We observe that LEO is biased towards generating “left” $97\%$ of the time. However, the ground-truth distribution of “left” and “right” should be balanced, suggesting that LEO may have a limited understanding of situated spatial relationships. In contrast, our model produces a distribution that closely matches the ground truth, demonstrating improved situated understanding.
Qualitative Examples.
In Fig. 7, we showcase several successful examples to demonstrate the effectiveness of Spartun3D-LLM across various situated tasks. Notably, in Fig 7 (c), the model without an explicit alignment module tends to generate more general or vague spatial descriptions, such as “turn around”. In contrast, with the alignment module, the model produces more specific details, including terms like “turn slightly right”. To verify this, we examine $30$ examples from both situated planning and situated captioning tasks and observe this phenomenon in $17$ of them. This highlights how the proposed spatial alignment module enhances the generation of fine-grained spatial information, leading to more precise and contextually accurate outputs.
@@ 6 Discussion and Conclusion
In this work, our goal is to address the limitation of situated understanding of the 3D-based LLMs from two perspectives. First, we propose a method to construct an LLM-generated dataset based on our designed situated scene graph. Then, we propose an explicit situated spatial alignment on the 3D-LLM to encourage the model to learn alignment between 3D object and their textual representations directly. Finally, we provide comprehensive experiments to show our own benchmark and improve situated understanding of SQA3D and navigation. We also provide analysis to show our proposed explicit alignment module helps spatial understanding.
@@ Appendix A Appendix Section
A.1 Evaluation Metrics
@@
  • EM: A stringent evaluation metric that checks whether the predicted output matches the reference exactly, often used in tasks like question answering.
Figure 8: Word distribution of questions of object attribute and relation.
A.2 Implementation Details.
@@
A.3 Different Text Encoder.
We conducted an ablation study to evaluate the impact of different text encoders on enhancing 3D visual representations. Table 7 shows the results on object attribute and relation tasks. Our results indicate that text encoders from VL models outperform pure text encoders, highlighting the importance of multimodal learning for 3D object understanding. Among the tested encoders, the CLIP encoder yield the best results.
@@
Table 8: Prompts for Different Tasks.
What Is How Can Which Others
CLIPBERT (Ma et al., 2022) $39.7$ $46.0$ $40.5$ $45.6$ $36.1$ $38.4$
SQA3D (Ma et al., 2022) $31.6$ $63.8$ $46.0$ $69.5$ $43.9$ $45.3$
3D-LLM (Hong et al., 2023) $35.0$ $66.0$ $47.0$ $69.0$ $\mathbf{48.0}$ $46.0$
3D-Vista (Zhu et al., 2023) $34.8$ $63.3$ $45.4$ $\mathbf{69.8}$ $47.2$ $48.5$
LEO (Huang et al., 2023) $46.8$ $64.1$ $47.0$ $60.8$ $44.2$ $54.3$
Figure 9: SQA3D Analysis.

### v2

@@ 5.2 Experimental Results
Spartun3D Benchmark. We evaluate the performance of both the LEO model and Spartun3D-LLM after fine-tuning them on our proposed Spartun3D dataset. The fine-tuned LEO model is referred to as LEO+Spartun3D. Table 2 and Table 4 show the experimental results on captioning and QA tasks, respectively. We experiment with two different LLM backbones: Opt1.3B and Vicuna7B. Our experiments show that Spartun3D-LLM consistently outperforms LEO+Spartun3D across all question types (around $2\%-3\%$ across all metrics), regardless of the LLM backbone used, indicating the effectiveness of our explicit alignment module. We observe that initializing our model with LEO pre-trained weights improves performance. Notably, without fine-tuning, LEO performs reasonably well on attribute and relation questions in a zero-shot setting but struggles with other situated tasks.
@@
SQA3D Performance. We evaluate our method on the SQA3D dataset, whose scenes are derived from ScanNet (Dai et al., 2017). Their scenes differ from those in Spartun3D, which are sourced from 3RScan. We experiment with two settings: zero-shot and fine-tuning. In the zero-shot setting, we re-trained LEO on their dataset (row#1) only constructed from 3RScan excluding all dataset constructed from ScanNet to ensure a fair comparison with our method. As shown in Table 3, LEO performs poorly on SQA3D in the zero-shot setting, suggesting its limitations in learning situated understanding from its dataset. In contrast, LEO trained on Spartun3D (row#2) shows significant improvement, demonstrating the effectiveness of our dataset. Further comparisons of Spartun3D-LLM with LEO+Spartun3D demonstrate a better zero-shot learning (i.e., generalization) capability of our model. In the fine-tuning setting, Spartun3D-LLM continues to outperform LEO across all metrics. Table 8 in the Appendix provides a breakdown of the fine-tuned performance across various question types, where Spartun3D-LLM shows consistent improvement.
@@
Navigation Performance. To demonstrate the effectiveness of our approach on downstream embodied tasks, we evaluate it on the object navigation tasks. Specifically, we randomly select $5$ scenes that contain around $1000$ examples from the MP3D ObjNav dataset. In this task, we additionally input 2D ego-centric images to both LEO and Spartun3D-LLM for comparison. There are four types of navigation actions: turn left, turn right, move forward, and stop. We evaluate whether the model generates correct action at each step. We conduct the experiment in a zero-shot setting, and Table 5 shows the accuracy of the model’s performance. The baseline model, LEO, struggles to generate the required action-related text to guide navigation steps without fine-tuning specifically for navigation tasks. In contrast, our model demonstrates strong transferability to generate correct actions. Fig 8 (e) showcases a qualitative example, illustrating how our model effectively generates accurate navigation actions without task-specific fine-tuning.
@@
Figure 8: Qualitative Examples. (a), (c), (c) situated qa examples of object attribute and relation, affordance, and planning, respectively. (d) situated captioning. (e) navigation in a zero-shot setting. Our model generates both actions and descriptions of surrounding objects while navigating.
@@ 5.3 Ablation Study and Extra Analysis
Explicit Alignment Enhances General Spatial Understanding. We evaluate the effectiveness of our proposed situated spatial alignment module on general scene understanding tasks, such as Scan2Cap (Chen et al., 2021) and ScanQA (Azuma et al., 2022). In line with our approach for situated tasks, we construct textual descriptions for each object based on its attributes and spatial relations to others from a top-view perspective. As shown in Tab. 6, by incorporating the explicit spatial alignment module, our model shows better results, indicating that our proposed alignment module not only improves situated understanding but also enhances general 3D scene understanding.
@@
Improved Situated Understanding. To analyze the model’s situated understanding ability further, we visualize the distribution of model responses generated for questions requiring strong spatial understanding from SQA3D. Specifically, we extract questions starting with “which direction”. Fig. 10 illustrates the distribution of generated “directions”, including “left”, “right”, “forward” and “backward”. We observe that LEO is biased towards generating “left” $97\%$ of the time. However, the ground-truth distribution of “left” and “right” should be balanced, suggesting that LEO may have a limited understanding of situated spatial relationships. The bias is significantly mitigated when LEO is trained on our dataset (LEO+Spartun3D). While adding our alignment loss (Spartun3D-LLM) helps futher, our dataset is the primary factor in addressing the bias.
Scaling Performance. We conduct scaling experiments to demonstrate how model performance improves with the addition of Spartun3D datasets. As shown in Fig. 10, we evaluate performance on SQA3D and observe consistent improvement as the dataset scales, highlighting the potential for dataset expansion using our proposed method.
Qualitative Examples. In Fig. 8, we showcase several successful examples to demonstrate the effectiveness of Spartun3D-LLM across various situated tasks. Notably, in Fig 8 (c), the model without an explicit alignment module tends to generate more general or vague spatial descriptions, such as “turn around”. In contrast, with the alignment module, the model produces more specific details, including terms like “turn slightly right”. To verify this, we examine $30$ examples from both situated planning and situated captioning tasks and observe this phenomenon in $17$ of them. This highlights how the proposed spatial alignment module enhances the generation of fine-grained spatial information, leading to more precise and contextually accurate outputs.
@@ 6 Discussion and Conclusion
In this work, our goal is to address the limitation of situated understanding of the 3D-based LLMs from two perspectives. First, we propose a method to construct an LLM-generated dataset based on our designed situated scene graph. Then, we propose an explicit situated spatial alignment on the 3D-LLM to encourage the model to learn alignment between 3D object and their textual representations directly. Finally, we provide comprehensive experiments to show our own benchmark improve situated understanding of SQA3D and navigation. We also provide analysis to show our proposed explicit alignment module helps spatial understanding.
7 Acknowledgement
This project is partially supported by the Office of Naval Research (ONR) grant N00014-23-1-2417, the award No. 2238940 from the Faculty Early Career Development Program (CAREER) of the National Science Foundation (NSF) and the U.S. DARPA FoundSci Program #HR00112490370.. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of Office of Naval Research.
@@ Appendix A Appendix Section
A.1 Human Evaluation
Human Scores.
To evaluate the quality of Spartun3D based on human scores, we mainly consider the following two aspects.
  • Language Naturalness evaluates the flow and syntax of the language, ensuring it reads naturally as if written by a human and adheres to common-sense knowledge in practical scenarios. Examples of different scoring levels for situations, questions, and answers are provided in Table 7. For instance, a score of 1 might be assigned to a scenario like “Standing beside a blanket while there is a toilet in the background,” which is uncommon in a typical setting. Similarly, questions such as “Which is cleaner, the trash bin or the bed?” are unlikely to be asked by humans, reflecting low language naturalness.
  • Spatial Fidelity is critical to ensure that the generated data accurately represents the information in 3D scenes and adheres to factual details. This includes verifying the presence of objects within the scene and ensuring that spatial relationships between objects are correctly represented. Additionally, common sense knowledge must be considered, especially when unusual objects or scenarios are mentioned in the questions or answers. For example, as shown in Fig. 12, a score of 1 is assigned to the instance where “clothes are hung on a mirror.” This error arises because the human-annotated data from 3RScan labeled the mirror’s affordance as “hanging,” which misled the GPT model into generating an incorrect dataset.
Error Analysis.
We randomly sampled 50 examples from each task (200 in total) to validate the quality of our automatically generated data and manually assess the quality from the aspects of language naturalness and spatial fidelity. Language naturalness evaluates whether the generated texts are natural or written by a human, while spatial fidelity ensures that data accurately reflects the 3D scene. We observe 26 errors in total and summarize them into the following categories.
  • Semantic Errors: The generated sentences may contain semantic mistakes. For instance, the answer “ You should go in front of you to water the potted plant.”.
  • Common Sense Violations: The generated content may occasionally conflict with basic common sense knowledge, such as producing unusual questions and answers. For example, it might generate a question like “If I want to store items, which object is most accessible?” with the answer being “trash bin.” This issue arises because the human-annotated data includes an affordance for trash bins as objects for storing items. Such annotations inadvertently influence GPT-4o to generate QA pairs that conflict with common sense knowledge.
  • Spatial Inconsistencies: Errors in capturing or reasoning about spatial relationships in the 3D environment primarily occur in Situated Planning tasks, which demand complex two-step spatial reasoning. These errors often arise because the second action depends on the outcome of the first action, and inaccuracies sometimes occur during the generation of the second action.
  • Misalignment between visual context and textual descriptions. In some cases, the agent’s view is obstructed due to the room layout or object size. For example, consider the situation: “Standing beside the sofa, there is a closet on your left.” However, the closet is actually located in another room and cannot be seen from the agent’s current standpoint. To address this issue, we designed a scenario where the agent stands beside a pivot object and consistently faces the center of the pivot object, rather than facing a random object that could potentially be obstructed. Additionally, we incorporated pass-by spatial information to enhance the agent’s awareness of surrounding objects, providing a more comprehensive sense of the environment.
Figure 9: Scaling Effect on SQA3D.
Figure 10: SQA3D Labels Distribution
A.2 Experimental Results
A.2.1 Evaluation Metrics
@@
  • EM: An evaluation metric that checks whether the predicted output matches the reference exactly, often used in tasks like question answering.
Scores Situation Question Answer
5 You are standing beside the nightstand while there is a picture on your left. How many sofa chairs are to my left? You can use the lamp behind you, but be careful of the stools you will pass by.
4 You are standing beside a lamp while there is a box on your left. Is the bag closer to me than the trash bin behind me? You can place items on the cabinet to your left, which is the closest option. Alternatively, you can place them on the bed to your left.
3 You are standing beside the radiator while there is a curtain rod on your back. I want to read a book. Should I walk to the window or sit in the chair? You can go to the clothing hanging on the door behind you.
2 You are standing beside salt while there is a kitchen cabinet on your left. What are the characteristics of the window to my left? Turn slightly back to your right and head towards the lamp to move it closer to your bed. You may pass by the curtain rod.
1 You are standing beside a blanket while there is a toilet on your back. Which is cleaner, the trash bin behind me or the bed to my left? You should go in front of you to water the potted plant.
Table 7: Examples to Evaluate Language Naturalness.
A.3 Implementation Details.
@@
What Is How Can Which Others
CLIPBERT (Ma et al., 2022) $39.7$ $46.0$ $40.5$ $45.6$ $36.1$ $38.4$
SQA3D (Ma et al., 2022) $31.6$ $63.8$ $46.0$ $69.5$ $43.9$ $45.3$
3D-LLM (Hong et al., 2023) $35.0$ $66.0$ $47.0$ $69.0$ $\mathbf{48.0}$ $46.0$
3D-Vista (Zhu et al., 2023) $34.8$ $63.3$ $45.4$ $\mathbf{69.8}$ $47.2$ $48.5$
LEO (Huang et al., 2023) $46.8$ $64.1$ $47.0$ $60.8$ $44.2$ $54.3$
Figure 11: Spartun3D Word Distribution.
Figure 12: Examples to Evaluate Spatial Fidelity.
@@
A.4 Extra Experimental Results
Different Text Encoder.
We conducted an ablation study to evaluate the impact of different text encoders on enhancing 3D visual representations. Table 9 shows the results on object attribute and relation tasks. Our results indicate that text encoders from VL models outperform pure text encoders, highlighting the importance of multimodal learning for 3D object understanding. Among the tested encoders, the CLIP encoder yields the best results.
Table 10: Prompts for Different Tasks.